# Diameter and mixing time of the giant component in the percolated hypercube

### v2

@@ 4. Upper tail
Recall the survival probability $y(c)$. It is a classic fact [Dwa69, Ott49, Pit97] that a Galton-Watson process with offspring distribution Poisson($c$) has order $k\geq 1$ with probability $\mathrm{e}^{-ck}(ck)^{k-1}/k!$, and is infinite with probability
@@ 5.2. Typical properties of $G_{1}$ and $G_{2}$
The idea that ‘the giant component is well-spread’ in the percolated hypercube has been used (sometimes implicitly) in many previous works [AKS81, BKL92, DEKK24, EKK22]. The first lemma in this subsection gives a quantitative characterisation of this spreadness.
@@ 6.2. Proof of Theorem 1(b)
First, we introduce some terminology related to Markov chains (for a comprehensive introduction, see [LPW17]). For a graph $G$ with vertex set $V$ and a vertex $u\in V$, the lazy (simple) random walk on $G$ starting at $u$ is a Markov chain with initial state $u$ and such that, at every step, the walk stays at its current state $v$ with probability $1/2$ and moves to a uniformly chosen neighbour of $v$ in $G$ with probability $1/(2\deg(v))$. If $G$ is a connected graph, the lazy random walk on $G$ is an irreducible, reversible and ergodic Markov chain with stationary distribution $\pi$ on $V$ where $\pi(v)=\deg(v)/(2e(G))$. Our goal is to determine the speed of convergence of the lazy random walk to its stationary distribution. To this end, we define the total variation distance $d_{\mathrm{TV}}$ between two distributions $\pi_{1}$ and $\pi_{2}$ on $V$ by setting
@@
and if no set $S$ with the required properties exists for some $\rho\in[\pi_{\min},1]$, we set $\Phi(\rho)=1$. The following theorem of Fountoulakis and Reed [FR07a] bounds the mixing time of the lazy random walk on $G$ through the conductance of connected sets.
Theorem 6.4 (Theorem 1 of [FR07a]).
@@
We move to part (b). It is enough to show that whp no set $S\subseteq V(L_{1})$ satisfies $\pi(V(L_{1})\setminus S)\geq 1/2$ and $|V(L_{1}\setminus S)|\leq K_{2}^{-1}v(L_{1})$. By [DKL25, Theorem 3.1], $G_{2}$ converges locally in probability to a Galton-Watson process with offspring distribution Poisson($c$); see the definition of the local convergence in probability in [DKL25]. Here, we will need the following simple corollary from [DKL25, Theorem 3.1]: for every integer $i\geq 0$, whp the number of vertices in $G_{2}$ with degree $i$ is at most $n\exp(-\Theta(\sqrt{\log d}))$-far from $n\cdot\mathbb{P}(\mathrm{Po}(c)=i)$.
@@ 7. Diameter
In [ADEK23], it was shown that if $p=\frac{c}{d}$ with $c>\mathrm{e}$, then with probability bounded away from zero, $\mathbf{0}$ is connected to $\mathbf{1}$ in a ’monotone’ path of length $d$. It was further shown that if $p=\frac{c}{d}$ with $c<e$, then whp this does not hold. The next result, which is the key technical lemma in this section, shows that even when $1<c\leq e$, there is a non-negligible probability of connecting $\mathbf{0}$ and $\mathbf{1}$ with a path of length $O(d)$. Its proof is delayed to Subsection 7.1.
@@
Since $pd=c>1$, whp the second-largest component in $Q^{d}_{p}$ has order $O(d)$, see e.g. [FK16, Theorem 13.1]. We assume this event in the sequel. Fix $v=\bf{0}$ and a sufficiently small $\delta>0$ satisfying $p(1-6\delta)d>1+\delta$. Let $J=\{1,\cdots,\delta d\}$ and let $W$ be a set consisting of $\delta^{2}d^{2}/4$ vertices with support of size 2 and included in $J$. In particular, all vertices in $W$ are within distance two from $v$ in $Q^{d}$. Further, for all $u,w\in W$, $Q(u,J)$ and $Q(w,J)$ are vertex-disjoint, and each has dimension $(1-\delta)d$.
@@
Classic martingale arguments for branching processes (see e.g. [D19, Chapter 4.3.4]) show that the BFS process reaches a layer $i\leq(\log d)^{2}$ with $z_{i}=\delta d$ with probability $\eta=\eta(\delta)>0$ which does not depend on $d$; denote this event by $\mathcal{G}_{u}^{\prime}$. Moreover, for every $i\leq\delta d-1$ with $z_{i}\geq\delta d$, Chernoff’s bound implies that

### v3

@@ 4. Upper tail
Recall the survival probability $y(c)$. It is a classic fact [24, 49, 51] that a Galton-Watson process with offspring distribution Poisson($c$) has order $k\geq 1$ with probability $\mathrm{e}^{-ck}(ck)^{k-1}/k!$, and is infinite with probability
@@ 5.2. Typical properties of $G_{1}$ and $G_{2}$
The idea that ‘the giant component is well-spread’ in the percolated hypercube has been used (sometimes implicitly) in many previous works [1, 13, 21, 25]. The first lemma in this subsection gives a quantitative characterisation of this spreadness.
@@ 6.2. Proof of Theorem 1(b)
First, we introduce some terminology related to Markov chains (for a comprehensive introduction, see [45]). For a graph $G$ with vertex set $V$ and a vertex $u\in V$, the lazy (simple) random walk on $G$ starting at $u$ is a Markov chain with initial state $u$ and such that, at every step, the walk stays at its current state $v$ with probability $1/2$ and moves to a uniformly chosen neighbour of $v$ in $G$ with probability $1/(2\deg(v))$. If $G$ is a connected graph, the lazy random walk on $G$ is an irreducible, reversible and ergodic Markov chain with stationary distribution $\pi$ on $V$ where $\pi(v)=\deg(v)/(2e(G))$. Our goal is to determine the speed of convergence of the lazy random walk to its stationary distribution. To this end, we define the total variation distance $d_{\mathrm{TV}}$ between two distributions $\pi_{1}$ and $\pi_{2}$ on $V$ by setting
@@
and if no set $S$ with the required properties exists for some $\rho\in[\pi_{\min},1]$, we set $\Phi(\rho)=1$. The following theorem of Fountoulakis and Reed [29] bounds the mixing time of the lazy random walk on $G$ through the conductance of connected sets.
Theorem 6.4 (Theorem 1 of [29]).
@@
We move to part (b). It is enough to show that whp no set $S\subseteq V(L_{1})$ satisfies $\pi(V(L_{1})\setminus S)\geq 1/2$ and $|V(L_{1}\setminus S)|\leq K_{2}^{-1}v(L_{1})$. By [22, Theorem 3.1], $G_{2}$ converges locally in probability to a Galton-Watson process with offspring distribution Poisson($c$); see the definition of the local convergence in probability in [22]. Here, we will need the following simple corollary from [22, Theorem 3.1]: for every integer $i\geq 0$, whp the number of vertices in $G_{2}$ with degree $i$ is at most $n\exp(-\Theta(\sqrt{\log d}))$-far from $n\cdot\mathbb{P}(\mathrm{Po}(c)=i)$.
@@ 7. Diameter
In [3], it was shown that if $p=\frac{c}{d}$ with $c>\mathrm{e}$, then with probability bounded away from zero, $\mathbf{0}$ is connected to $\mathbf{1}$ in a ’monotone’ path of length $d$. It was further shown that if $p=\frac{c}{d}$ with $c<e$, then whp this does not hold. The next result, which is the key technical lemma in this section, shows that even when $1<c\leq e$, there is a non-negligible probability of connecting $\mathbf{0}$ and $\mathbf{1}$ with a path of length $O(d)$. Its proof is delayed to Subsection 7.1.
@@
Since $pd=c>1$, whp the second-largest component in $Q^{d}_{p}$ has order $O(d)$, see e.g. [31, Theorem 13.1]. We assume this event in the sequel. Fix $v=\bf{0}$ and a sufficiently small $\delta>0$ satisfying $p(1-6\delta)d>1+\delta$. Let $J=\{1,\cdots,\delta d\}$ and let $W$ be a set consisting of $\delta^{2}d^{2}/4$ vertices with support of size 2 and included in $J$. In particular, all vertices in $W$ are within distance two from $v$ in $Q^{d}$. Further, for all $u,w\in W$, $Q(u,J)$ and $Q(w,J)$ are vertex-disjoint, and each has dimension $(1-\delta)d$.
@@
Classic martingale arguments for branching processes (see e.g. [23, Chapter 4.3.4]) show that the BFS process reaches a layer $i\leq(\log d)^{2}$ with $z_{i}=\delta d$ with probability $\eta=\eta(\delta)>0$ which does not depend on $d$; denote this event by $\mathcal{G}_{u}^{\prime}$. Moreover, for every $i\leq\delta d-1$ with $z_{i}\geq\delta d$, Chernoff’s bound implies that